You work for a fencing company. A customer called and wanted to fence in his 1,320 square- foot garden. He ordered 148 feet of fencing, but you forgot to ask him for the width and length of the garden. You don't want to bug him again. what is the length and width of the garden? what are the dimensions?
Answer by lwsshak3(11628)

You work for a fencing company. A customer called and wanted to fence in his 1,320 square- foot garden. He ordered 148 feet of fencing, but you forgot to ask him for the width and length of the garden. You don't want to bug him again. what is the length and width of the garden? what are the dimensions?
**
let L=length
let W=width
2L+2W=148
L+W=74
W=74-L
L *W=area=1320
L(74-L)=1320
74L-L^2=1320
L^2-74L+1320=0
(L+30)(L+44)=0
L=30
or
L=44
ans:
Dimensions:
Length=44 ft
Width=30 ft